high pressure steam enters a turbine with a velocity of 125 ft/s and leaves the exhaust trunk of the turbine with a velocity of 1000 ft/s. calculate the kinetic energy in ft-lbf/lb and BTU/ lb at a. at the entrance to the turbine, and b. the exhaust trunk
1
Expert's answer
2020-02-28T10:22:36-0500


125 ft/s=38.1 m/s

1000 ft/s=304.8 m/s


"P_1=\frac{KE_1}{m}=\frac{v_1^2}{2}=\frac{38.1^2}{2}\approx726J/kg=0.726kJ/kg"


"P_2=\frac{KE_2}{m}=\frac{v_2^2}{2}=\frac{304.8^2}{2}\approx92903J/kg=92.903kJ/kg"


(a) 0.726 (kJ/kg)=242.73 (ft-lbf/lb)=0.312BTU/lb


(b) 92.903 (kJ/kg)=31073.44 (ft-lbf/lb)=39.941 (BTU/lb)